Grade 12th passMechanics

Why does the acceleration of a freely falling object not depend on the weight of the object?

Salman khan
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear Salman
 
Force acting on a free falling object (F) from rest is equal to the mass of the object (m) times acceleration of the free falling object (a).So, F = ma …… (1)But at the same time when the object is released from rest near the surface of earth, the down ward force of the Earth’s gravity acting on the object is called the weight of the object (W). Weight of the free falling object is equal to the mass of the object times (m) acceleration due gravity on that surface (g).Therefore, W = mg …… (2)From equation (1), free falling acceleration will be,a = F/m …… (3)The weight of the object (W) is equal to the force acting on the free falling object (F).So, W = F …… (4)Putting the value of equation (4) in equation (3), we get,a = F/m= W/m= mg/ m (Since, W = mg)= g …… (5)Thus from equation (5) we conclude that, the acceleration of a free falling object does not depend on the weight of the object.
 
Regards
Arun (askIITians forum expert)